# Out-of-Distribution Detection in Molecular Complexes via Diffusion Models for Irregular Graphs

### v1

@@ 4 Discussion
This work aims to contribute meaningfully to the debate between generalization and memorization of machine learning models. While new models frequently report strong generalization, such claims are often accompanied by concerns that performance may largely stem from memorizing recurring patterns shared between training and test data. This concern is particularly relevant for models trained on complex data such as protein-ligand interactions, where structural homologies are non-trivial to detect and identical interaction patterns can occur in otherwise dissimilar complexes. We argue that potential data leakage and distributional overlap should be analyzed at the level of the data representation accessible to the model, rather than at the raw structural data. This way can we rigorously evaluate whether training and test samples share similar representations for the model to exploit.
By explicitly relating prediction errors to OOD scores, we can disentangle generalization from dataset-specific memorization. The steepness of the increase in prediction errors with higher log-likelihood can serve as a quantitative marker of generalization ability. The rate at which prediction errors increases as a function of the log-likelihood can serve as a quantitative proxy for generalization capability. A steep increase implies that model performance relies on distributional overlap, degrading rapidly for OOD data. Conversely, a shallow increase indicates robust generalization, demonstrating the model’s resilience in regions of chemical and structural space that are less-represented within its training data. This enables direct comparisons across competing models, moving beyond single benchmark values. In this sense, OOD quantification can act as a certificate for benchmark datasets. Comparing the distribution of OOD log-likelihoods with respect to the training data clarifies whether high benchmark performance reflects actual generalization or mere memorization.
@@
We emphasize that our method should be read as one concrete instantiation of a broader and robust OOD detection blueprint. The core requirement is simply a generative dynamical model–a diffusion or flow-matching model–trained to represent the in-distribution of irregular graphs. Once such a model is available, the induced probability-flow dynamics provide a natural coordinate system in which to build trajectory-level diagnostics: one can design score functions and summary statistics that probe whichever structural regularities are relevant for the application, and then detect OOD by identifying trajectories whose behavior is atypical under the training distribution.
For instance, in protein-ligand settings one can augment likelihoods with diffusion-time features that explicitly track (i) geometric stability (e.g. preservation of pocket shape, RMSD-like drift of key residues, local rigidity/strain, or deviations from expected bond-length/angle priors after bond inference), (ii) interaction physics (e.g. evolution of hydrogen-bond patterns, salt-bridge and $\pi$-stacking persistence, steric clashes, packing density, solvent-exposed surface area proxies, or Coulomb/Lennard-Jones-style surrogate energies computed from coordinates), and (iii) biological coupling signals (e.g. co-folding/co-adaptation descriptors such as correlated motions of residue substructures with ligand substructures, contact-map consistency over diffusion time, or stability of secondary-structure–like local motifs in the pocket neighborhood). More broadly, any measurable property of the evolving state–including constraints, symmetries, conservation-like quantities, or domain-specific interaction scores–can be turned into a trajectory statistic and combined into a multivariate typicality model, yielding an OOD detector tailored to the scientific failure modes of interest.
While this work provides a proof of concept demonstrating that diffusion PF-ODE trajectory features enable successful OOD quantification, significant potential remains to improve the diffusion model’s sensitivity to chemical differences between ligands. Among the nine datasets tested, we identified one OOD dataset of metalloenzymes (carbonic anhydrases) that the model consistently flagged as in-distribution. We hypothesize that the current model struggles to recognize unique chemistry embedded in common scaffold structures. While the ligands in these complexes possessed simple, common scaffolds structures, they were characterized by high sulfur content and the presence of sulfonamide groups interacting with a zinc ion. This kind of chemistry is highly underrepresented in the dataset. Future work should focus on improving the model’s chemical understanding to better capture such distinct chemistry within otherwise familiar geometries.
@@ 5.1 Datasets
This dataset split was employed for both the diffusion model and the GEMS binding affinity prediction model. The diffusion model was trained on the training dataset. Validation losses were computed throughout the training process using the validation dataset. All non-training datasets (validation, CASF2016, and OOD sets) were subsequently used to evaluate the model’s ability to differentiate and rank datasets according to their OOD level relative to the training data. GEMS was trained exclusively on the training dataset using 5-fold cross-validation (CV). All other datasets served as test sets to evaluate generalization capability to datasets of varying distribution shifts, ranging from IID (e.g. the validation set) to highly OOD (e.g. the HIV proteases).
@@ 5.2.1 Background and Motivation
Continuous diffusion and probability–flow.
@@
[…]
@@
[…]
@@
Denoising as deconvolution (posterior mean).
@@ 5.2.2 Unified continuous diffusion for molecular coordinates and identities
Background: score interpolation for categorical data.
@@
Learning molecular embeddings without collapse.
@@ 5.2.3 Diffusion Model Setup
Divergence via Hutchinson under COM constraints.
@@ 5.2.4 Diffusion Model Training and Evaluation
Ligand atom coordinates and types were extracted from the Gemmi objects (hydrogens omitted). Ligand atom coordinates were combined into a ligand feature matrix ($\mathbf{C}_{L}$) and ligand atom types were one-hot-encoded to construct the ligand feature matrix ($\mathbf{F}_{L}$) using the following 10 categories:
@@
where $N_{L}$ is the number of ligand atoms and $N_{P}$ is the number of pocket residues. These four arrays were then combined in a PyTorch Geometric Data object for each complex. To construct datasets, these Data objects were combined into PyTorch datasets according to the splitting outlined in section 5.1.
@@
To maximize the model’s performance and ensure robust training, we performed a systematic hyperparameter optimization. The search space was defined across several key architectural and training features, including learning rate, the number of sampling steps, the number of EGNN layers, the joint and hidden embedding sizes and the edge feature dimensionality. Initially, we employed a randomized search with short training runs to set reasonable search space boundaries. Then we explored the resulting search space outlined in Table 3 using Bayesian optimization with the Optuna (v.4.4.0) python package. The best parameter combination was selected based on a joint assessment of validation loss and sampling-based evaluations, including percent of fragmented ligands, mean number of ligand fragments, mean number of rings, mean ring size and Jensen-Shannon divergences between training and sampling ligand atom and protein residue type distributions. The model was trained using the final configuration (in the rightmost column of the Table 3), resulting in a model with with 6’386’128 parameters. The training process was continued for as long as the training and validation losses continued to decrease, but was stopped eventually after 1400 epochs. The last saved model (epoch 1390) was used for log-likelihood based OOD evaluation. Protein-ligand interactions sampled from this model contained ligands with 92% chemical validity, 72% of unfragmented ligands and a mean number of 2.34 rings with mean size of 5.31 per molecule. Divergences in atom and amino acid type distributions (Jensen–Shannon) from the training data were very low, at 0.0017 and 0.004, respectively.
Table 3: Hyperparameter search spaces and selected optimal values.
@@ 5.3 GEMS Training and Evaluation
Custom datasets were constructed for the training, validation, CASF2016, and OOD sets. This involved computing graph representations of the complexes across the entire PDBbind v.2020 dataset, featurized using language model embeddings derived from ESM2 [17], Ankh [18], and ChemBERTa2 [19].
The GEMS model was trained on the training dataset using 5-fold cross-validation (CV) and the default hyperparameters:
@@
The model was trained using early stopping, terminating the process when the validation RMSE did not decrease for 100 epochs.
@@
The coefficient of determination ($R^{2}$) was chosen as the main readout for model performance. Other common regression metrics, such as the Pearson Correlation Coefficient (PCC) and Root Mean Squared Error (RMSE), were disregarded due to specific limitations. PCC is insensitive to scale offsets, meaning a high correlation score can coexist with large absolute errors. Conversely, prediction RMSE values are sensitive to variations in the label scale and variance across different datasets, distorting performance comparisons. For datasets with a narrow label distribution, naive models (which predict the dataset mean for all complexes) can achieve deceptively low RMSE values, whereas the same naive models would achieve poor RMSE on datasets with a wide label distribution. $R^{2}$ normalizes performance relative to the target variance, providing a fairer comparison across diverse datasets. To compute complex-wise errors, we applied a similar normalization to account for varying label spread across datasets. Specifically, we used Variance-Normalized Errors (VNE), calculated by dividing the absolute error of each prediction by the variance of the dataset’s labels. This effectively rescales the error relative to the difficulty of the dataset, providing a metric that reflects how the model improves sample scoring over a naive baseline that simply predicts the dataset mean.
@@ 5.4 Protein-Ligand Complex Similarity Computation
To compute a similarity score between two protein-ligand complexes, Tanimoto ligand similarity, protein similarity (TM scores), and a pocket-aligned ligand root-mean-square-deviation were combined into an overall structure-based similarity score. This combination captures the presence of similar interaction patterns between two protein-ligand complexes on a structural level. These similarity scores were precomputed in our earlier work [3] and were reused here:
@@ 5.5 Trajectory-aware out-of-distribution detection
Feature construction.
We aggregate d=19d{=}19 PF-ODE trajectory statistics per complex (full list in Fig. S2). For exposition we display only the most discriminative coordinates:
@@
Class-conditional density model.
@@
Scoring and decision rule (test time).
@@
Bootstrap evaluation under class imbalance.
@@
Confidence and risk stratification.
@@
Why trajectory features help.
@@ 5.6 Embedding–space baseline on GEMS
Unsupervised detectors on ID embeddings.
@@
Calibration and decision rules.
@@ 5.7 Rate-In baseline on GEMS
Information-preserving dropout.
@@
Per-layer rate optimisation.
@@
OOD scoring from variance and optimisation dynamics.
@@ 5.9 Acknowledgments
S1.1 Trajectory-aware OOD detection successful for 3dd0 dataset
For the dataset of alpha-carbonic anhydrases 3dd0, the PF-ODE yielded unexpectedly high log-likelihoods, which would lead to a misclassification as in-distribution (Figure S4). While log-likelihoods alone failed to identify 3dd0 complexes as OOD, we extracted additional PF-ODE trajectory statistics that capture the diffusion model’s struggle to process OOD data. In this augmented feature space, our OOD detection model can draw a clear separation between training and 3dd0 samples and successfully corrected the prediction, recognizing 3dd0 as OOD with high confidence (Table S1).
OOD detection AUROC Acc F1 Spec Prec Recall
[…]
GEMS Rate-In $0.337\,{\pm}\,0.029$ $0.498\,{\pm}\,0.005$ $0.238\,{\pm}\,0.317$ $0.641\,{\pm}\,0.478$ $0.179\,{\pm}\,0.239$ $0.354\,{\pm}\,0.472$
GEMS Embedding $0.728\,{\pm}\,0.019$ $0.740\,{\pm}\,0.012$ $0.764\,{\pm}\,0.008$ $0.636\,{\pm}\,0.024$ $0.699\,{\pm}\,0.014$ $0.844\,{\pm}\,0.000$
Table S1: Trajectory-aware OOD detector identifies 3dd0 as OOD: Comparison of AUROC, accuracy (Acc), F1-Score, specificity (Spec), precision (Prec) and recall with mean $\pm$ standard deviation across OOD datasets for the trajectory-aware LDR diffusion-based detector (Traj-LDR), the embedding-space GEMS baseline, and the Rate-In baseline. The trajectory-aware OOD detector classifies the dataset of alpha-carbonic anhydrases (3dd0) as out-of-distribution with high confidence. In contrast, the GEMS-based baselines show reduced separation capability.
@@
Figure S1: Error-log-likelihood relation follows a scaled a shifted exponential: Scatterplot illustrating the relationship between the log-likelihoods assigned by the diffusion model (x-axis) and the corresponding GEMS binding affinity prediction errors (y-axis) across the validation, CASF2016, and the out-of-distribution (OOD) datasets. A hierarchical fitting of five scaled and shifted exponential curves shows that log-likelihoods provide a rough estimate of the expected GEMS prediction error. Across all non-training complexes ($N=6223$), 75.8% had predicted errors that fell within the bounds of the fitted exponential curves, with 8.84% high outliers (errors above the fit) and 15.39% low outliers (errors below the fit). To allow fair comparison across datasets with different label spreads, absolute errors are normalized by the label variance, preventing artificially low error values on narrowly distributed datasets. Exponential curves are fit to a maximum of 500 randomly sampled points from each non-training distribution. Only 100 randomly sampled complexes from each datasets are shown in the to improve visibility.
S1.3 Trajectory feature importance
To understand which aspects of the PF-ODE trajectory drive separation between in-distribution (ID) and out-of-distribution (OOD) complexes, we quantify the importance of each of our $d=19$ trajectory statistics and rank them accordingly. Rather than reporting a single heuristic, we use a combined importance score that rewards features which (i) separate ID and OOD strongly and (ii) do so consistently across complexes and bootstrap splits. Concretely, for each feature we compute an effect size between ID and OOD summaries (e.g. Cohen’s $d$) and a stability term (e.g. variance ratio / robustness under resampling), standardize these quantities across features, and aggregate them (additively) into one score.
00.20.20.40.40.60.60.80.8111.21.21.41.41.61.6log_likelihoodpath_tortuositypath_efficiencyvf_l2_stdvf_l2_maxvf_l2_meanmean_accelerationmean_lipschitzmax_lipschitzcoupling_consistencytotal_flow_energyvf_spikinesstotal_angular_deviationmax_intermol_changesmoothness_scorecoord_feature_ratiodynamic_coord_feature_couplingmean_com_driftmax_com_driftCombined importance score
Figure S2: Top PF-ODE trajectory features ranked by a combined importance score (separation strength + stability under resampling).
Several patterns stand out. First, PF-ODE log-likelihood is the single most discriminative coordinate, but it is not the whole story: multiple trajectory descriptors achieve comparable importance, confirming that OOD shift manifests as a dynamic signature along the probability flow rather than only as a terminal density value. Second, the highest-ranked non-likelihood features cluster into interpretable families: (i) path geometry (tortuosity/efficiency), indicating that OOD inputs induce longer and less direct transport to noise; (ii) field magnitude and burstiness (vf_l2 mean/max/std and spikiness), indicating larger and more intermittent corrective updates; and (iii) stability and smoothness (Lipschitz and acceleration), indicating stiffer and less temporally regular dynamics. Finally, coupling-related terms (e.g. coupling_consistency and dynamic_coord_feature_coupling) rank non-trivially, suggesting that OOD behavior is not purely geometric: atypical complexes also perturb how coordinate and identity updates co-evolve across noise scales.
S1.4 Evaluation of sensitivity to ligand size and chemical features
@@
A notable limitation of our diffusion model is the limited sensitivity to unique chemistry. We observed that the model struggles to detect distribution shifts when unique chemistry is masked by common geometries. In this context, the carbonic anhydrase dataset 3dd0 presented a distinct challenge. This cluster of metalloenzymes was consistently assigned high log-likelihoods (Figure S4) despite containing chemical features that are underrepresented in the training distribution, such as zinc ions and sulfonamide-rich ligands. Visual inspection reveals that these ligands often possess simple, generic scaffolds. This implies that the current diffusion model may overemphasize geometric familiarity at the expense of recognizing specific, rare chemical substructures.
Figure S4: OOD dataset 3dd0 yields high log-likelihoods: Distributions of log-likelihoods assigned to protein-ligand complexes belonging to the a) training and b) 3dd0 dataset of alpha carbonic anhydrases. Lower values signify increased deviation from the learned distribution (OOD) and higher values indicate high in-distribution (ID) probability. Means, medians, standard deviations (Std) and number of samples (N) are depicted on the right side of the histograms. Vertical dashed lines indicate the median of the respective log-likelihood distributions. The training distribution was randomly subsampled, showing only a fraction of all 10’510 complexes.
Figure S5: Distributions of train-test aggregated similarity scores: A composite score S=max(Tanimoto+TMScore+(1RMSD),0)S=max(Tanimoto+TMScore+(1-RMSD),0) was calculated as the sum of Tanimoto similarity, TM-scores, and inverted pocket-aligned ligand root mean squared deviation (RMSD). High scores signify highly similar protein-ligand complexes (3.0 = identical complexes). Each box represents the distribution of $N=10,510$ similarity scores. Specifically, these are the scores between each training dataset complex and its most similar counterpart in the respective test dataset, as measured by that specific metric. Boxplots show the median (centre line), 25th–75th percentiles (box), whiskers extend to data points within 1.5 × IQR, outliers are not shown.
S1.5 Intra-Cluster Ligand Diversity
@@
Figure S6: Intra-cluster ligand diversity of all datasets sorted by cluster tightness in ascending order. Three robust intra-cluster diversity metrics are included: (1 median distance to medoid measures typical cluster tightness by computing the median distance from the most representative point to all other points; (2 90th percentile radius captures robust cluster size by identifying the distance that encompasses 90% of points from the medoid; and (3 pairwise inter-quartile mean (IQM) quantifies overall cluster cohesion by averaging distances between all unique pairs within the middle 50% of the pairwise distance distribution, providing resistance to outliers.
S1.6 Sampling Evaluation
@@
S1.7 PF-ODE likelihoods as predictors of predictive error
@@
S1.7.1 Relative entropy along the PF-ODE
@@
KL divergence and relative Fisher information.
@@
[…]
@@
[…]

### v2

@@ 3 Discussion
We have introduced a generative framework for detecting out-of-distribution (OOD) structures in irregular 3D graph data, demonstrating that a diffusion model trained purely on raw geometry and discrete biochemical identities recovers a principled, task-agnostic OOD signal within its dynamic noising trajectories. Our methodology employs a unified continuous diffusion process and posterior-mean interpolation to establish a stable SE(3)-equivariant denoiser whose probability-flow ODE (PF-ODE) trajectories provide a per-complex notion of typicality with respect to the training distribution. The central methodological advancement is the augmentation of scalar likelihoods with 18 additional trajectory statistics. By capturing path geometry, flow stiffness, vector-field instability, and energetic cost, we demonstrate that OOD structures induce distinct dynamics in their diffusion trajectories that facilitate robust detection, effectively bypassing the complexity bias that often confounds methods based on likelihoods alone. This approach successfully identifies bioinformatically validated OOD protein-ligand interactions and outperforms classical post-hoc baselines despite working in a completely unsupervised setting.
The practical utility of this unsupervised approach is underscored by the massive disparity between available structural data and curated labels. With the amount of structural data growing at an exponential rate relative to labeled datasets like PDBbind, the ability to derive OOD signals without supervision is critical. To our knowledge, this work represents the first report of unsupervised OOD detection specifically for 3D geometric graph data, a modality where detection via simple statistical checks or similarity computations is inherently impractical due to the complex combination of topology, 3D geometry, and discrete features. Instead, effective detection requires a joint assessment of these features; a task our framework accomplishes by explicitly modeling the probability density of this joint distribution. Our results demonstrate that these diffusion-based estimates can provide an a priori assessment of a model’s reliability, essentially forecasting prediction errors before a complex is processed by a downstream predictor. This provides a critical safety measure for deployed machine learning models, warning users when input data points are atypical and the model’s output may be unreliable.
Beyond practical safety, this work contributes to the ongoing debate between generalization and memorization in structural bioinformatics. While modern models often report high benchmark performance, such claims are frequently accompanied by concerns that performance may largely stem from memorizing recurring patterns shared between training and test data. We propose that OOD quantification can act as a ”certificate” for benchmark datasets and the connected generalization claims. By explicitly relating prediction errors to trajectory-based OOD scores, we can disentangle true generalization from dataset-specific memorization. A model exhibiting a shallow increase in error as it moves into OOD territory demonstrates genuine generalization to data underrepresented in its training set, whereas a steep increase indicates performance rooted in memorized training motifs. This enables a direct comparison across competing architectures, moving beyond single-point benchmark values to evaluate a model’s true generalization capacity.
@@
While this work provides a proof of concept for PF-ODE trajectory features, significant potential remains to improve the diffusion model’s distribution learning to further increase sensitivity to subtle chemical and structural nuances. Nevertheless, our findings demonstrate that the generative modeling of 3D graphs inherently captures the ”essence” of the data distribution. The emergence of robust OOD signals as a byproduct of the generative learning process, without task-specific labels or hand-crafted priors, offers a versatile mechanism for safety and reliability in geometric deep learning. This framework is not restricted to molecular complexes but is inherently generalizable to any modality amenable to continuous diffusion, offering a general recipe for OOD quantification and error forecasting. Our method should be viewed as one concrete instantiation of a broader OOD detection blueprint: The core requirement is simply a generative dynamical model, a diffusion or flow-matching model, trained to represent the in-distribution. Once such a model is available, the induced probability-flow dynamics provide a natural coordinate system in which to build trajectory-level diagnostics. Researchers can design score functions and summary statistics that probe the structural regularities relevant to their specific application, detecting OOD samples by identifying trajectories whose behavior is atypical under the learned training distribution.
@@ 4.1 Datasets
This dataset split was employed for both the diffusion model and the GEMS binding affinity prediction model. The diffusion model was trained on the training dataset, with validation losses computed throughout the training process using the validation dataset. All non-training datasets (validation, CASF2016, and OOD sets) were subsequently used to evaluate the model’s ability to differentiate and rank datasets according to their OOD level relative to the training data. GEMS was trained exclusively on the training dataset using 5-fold cross-validation (CV). All other datasets served as test sets to evaluate generalization capability to datasets of varying distribution shifts, ranging from ID (e.g. the validation set) to highly OOD (e.g. the HIV proteases).
@@ 4.2.1 Background and Motivation
Continuous diffusion and probability–flow
@@
$\mathrm{d}\bm{x}\;=\;f(\bm{x},t)\,\mathrm{d}t\;+\;g(t)\,\mathrm{d}\bm{w},\qquad\bm{w}\text{ a standard Wiener process},$ (4.1)
@@
$\log p(\bm{x}_{0})\;=\;\log p_{1}(\bm{x}_{1})\;+\;\int_{0}^{1}\!\mathrm{tr}\!\left(\frac{\partial\tilde{f}}{\partial\bm{x}}(\bm{x}(t),t)\right)\mathrm{d}t,$ (4.3)
@@
Denoising as deconvolution (posterior mean)
@@ 4.2.2 Unified continuous diffusion for molecular coordinates and identities
Background: Score interpolation for categorical data
@@
Learning molecular embeddings without collapse
@@ 4.2.3 Diffusion Model Setup
Divergence via Hutchinson under COM constraints
@@ 4.2.4 Diffusion Model Training and Evaluation
Ligand atom coordinates and types were extracted from the Gemmi objects (hydrogens omitted). Ligand atom coordinates were combined into a ligand coordinate matrix ($\mathbf{C}_{L}$) and ligand atom types were one-hot-encoded to construct the ligand feature matrix ($\mathbf{F}_{L}$) using the following 10 categories:
@@
where $N_{L}$ is the number of ligand atoms and $N_{P}$ is the number of pocket residues. These four arrays were then merged in a PyTorch geometric data object for each complex. To construct datasets, these data objects were combined into PyTorch datasets according to the splitting outlined in section 4.1.
@@
To maximize the model’s performance and ensure robust training, we performed a systematic hyperparameter optimization. The search space was defined across several key architectural and training features, including learning rate, the number of sampling steps, the number of EGNN layers, the joint and hidden embedding sizes and the edge feature dimensionality. Initially, we employed a randomized search with short training runs to set reasonable search space boundaries. Then we explored the resulting search space outlined in Table 3 using Bayesian optimization with the Optuna (v.4.4.0) python package. The best parameter combination was selected based on a joint assessment of validation loss and sampling-based evaluations, including percent of fragmented ligands, mean number of ligand fragments, mean number of rings, mean ring size and Jensen-Shannon divergences between training and sampled ligand atom and protein residue type distributions. The model was trained using the final configuration (in the rightmost column of the Table 3), resulting in a model with with 6’386’128 parameters. The training process was continued for as long as the training and validation losses continued to decrease, but was stopped eventually after 1400 epochs. The last saved model (epoch 1390) was used for PF-ODE trajectory-based OOD evaluation. Protein-ligand interactions sampled from this model contained ligands with 92% chemical validity, 72% of unfragmented ligands and a mean number of 2.34 rings with mean size of 5.31 per molecule. Divergences in atom and amino acid type distributions (Jensen–Shannon) from the training data were very low, at 0.0017 and 0.004, respectively.
@@
Table 3: Hyperparameter search spaces and selected optimal values used for diffusion model training.
@@ 4.3 GEMS Training and Evaluation
Custom datasets were constructed for the training, validation, CASF2016, and OOD sets. This involved computing graph representations of the complexes across the entire PDBbind v.2020 dataset, featurized using language model embeddings derived from ESM2 [23], Ankh [9], and ChemBERTa2 [1]. The GEMS model was trained on the training dataset using 5-fold cross-validation (CV) and the default hyperparameters:
@@
The model was trained using early stopping, terminating the process when the validation RMSE did not decrease for 100 epochs. The last model that improved validation RMSE was then used.
@@
The coefficient of determination ($R^{2}$) was chosen as the main readout for model performance. Other common regression metrics, such as the Pearson Correlation Coefficient (PCC) and Root Mean Squared Error (RMSE), were disregarded due to specific limitations. PCC is insensitive to scale offsets, meaning a high correlation score can coexist with large absolute errors. Conversely, prediction RMSE values are sensitive to variations in the label spread across different datasets, distorting performance comparisons. For datasets with a narrow label distribution, naive models (which predict the dataset mean for all complexes) can achieve deceptively low RMSE values, whereas the same naive models would achieve poor RMSE on datasets with a wide label distribution. $R^{2}$ normalizes performance relative to the target variance, providing a fairer comparison across diverse datasets. To compute complex-wise errors, we applied a similar normalization to account for varying label spread across datasets. Specifically, we used Variance-Normalized Errors (VNE), calculated by dividing the absolute error of each prediction by the variance of the dataset’s labels. This effectively rescales the error relative to the difficulty of the dataset, providing a metric that reflects how much the model improves sample scoring over a naive baseline that simply predicts the dataset mean.
@@ 4.4 Protein-Ligand Complex Similarity Computation
To compute a similarity score between two protein-ligand complexes, ligand similarity (Tanimoto), protein similarity (TM scores), and a pocket-aligned ligand root-mean-square-deviation (RMSD) were combined into an overall structure-based similarity score. This combination captures the presence of similar interaction patterns between two protein-ligand complexes on a structural level. These similarity scores were precomputed in our earlier work [13] and were reused here:
@@ 4.5 Trajectory-aware out-of-distribution detection
Feature construction
We aggregate d=19d{=}19 PF-ODE trajectory statistics per complex (full list in Fig. 7). For exposition we display only the most discriminative coordinates:
@@
Class-conditional density model
@@
Scoring and decision rule (test time)
@@
Bootstrap evaluation under class imbalance
@@
Confidence and risk stratification
@@
Why trajectory features help
@@ 4.6 Embedding–space baseline on GEMS
Unsupervised detectors on ID embeddings
@@
Calibration and decision rules
@@ 4.7 Rate-In baseline on GEMS
Information-preserving dropout
@@
Per-layer rate optimisation
@@
OOD scoring from variance and optimisation dynamics
@@ 4.9 Acknowledgments
S1.1 Related Work
The problem of detecting out-of-distribution (OOD) samples has been extensively studied in computer vision. Early discriminative approaches relied on post-hoc analysis of supervised classifier outputs, such as the maximum softmax probability [15] or temperature scaling [22]. More recent methods leverage energy-based scores [24] or activation pruning [34] to separate ID from OOD data. However, these methods usually require training labels of the in-distribution data and their applicability is thus limited to labeled datasets.
Deep generative models for OOD detection
Generative models are particularly well suited for OOD detection in an unsupervised setting, as they are trained on the generative task and require no labels. During training, these models learn a probability density function of the training data. This allows to estimate the likelihood of a given new sample under the learned model [32, 12].
Some generative approaches for OOD detection are reconstruction-based, predominantly using Variational Autoencoders (VAEs) [2, 37, 3]. The underlying assumption is that a model trained on in-distribution data will fail to accurately reconstruct OOD samples, yielding a high reconstruction error. However, some research highlighted that standard autoencoders often reconstruct outliers perfectly, rendering the reconstruction error unreliable [5, 41].
To overcome this, research shifted toward explicit density estimation using the likelihood $p(x)$ directly to identify OOD samples. These methods model the in-distribution data with a probabilistic model and flag test data that lies in low-density regions as OOD. While theoretically sound, this direction faced a major setback when [27] demonstrated the likelihood paradox, where generative models assigned higher likelihoods to simple OOD data (e.g., SVHN) than to the complex training distribution (e.g., CIFAR-10). This was confirmed by following research, showing that likelihood exhibits a strong bias towards the input complexity of the data [30, 19]. Some works have attempted to mitigate this using likelihood ratios [28, 12], complexity-based corrections [30] or applying diffusion in the representation space of a pre-trained image encoder [7].
Deviating from pure likelihood-based methods, Heng et al. (2024) [16] introduced diffusion trajectory features, specifically the rate-of-change and curvature of the PF-ODE, as effective OOD signals for image benchmarks. Their trajectory-focused approach proved resistant to the complexity bias that often confounds likelihood-based OOD classification. However, while their study was limited to image data and a small set of descriptors, we translate these principles to the challenging domain of complex 3D graphs. We expand the feature space to 18 distinct trajectory features and model them jointly with log-likelihoods. This composite approach yields a robust OOD detector for 3D graph data, where likelihoods and trajectory features provide complementary, essential signals for classification in this domain.
OOD detection on graphs
Extending OOD detection from image data to graph domains introduces unique challenges due to the irregular, non-Euclidean nature of graph data. While several works have adapted discriminative scores for Graph Neural Networks (GNNs) trained with standard supervised objectives [36], unsupervised generative approaches typically focus on reconstructing the adjacency matrix conditioned on node features ($P(A|X)$). These methods operate on the assumption that a failure to recreate a graph’s structure indicates an outlier [21, 39]. However, existing unsupervised graph-level OOD detectors [26, 21, 31] generally operate on 2D topological graphs, neglecting the precise 3D geometry that is essential for molecular complexes. Shen et al. (2024) [31] use a diffusion model for OOD detection on molecular graphs of drug-like compounds. However, they also limit their modeling to topological graphs without 3D geometry, rendering their method unsuitable for our use case of OOD detection in protein-ligand interactions. In molecular interaction graphs, OOD characteristics manifest not only in features or connectivity but crucially in geometry. Since the distinction between in-distribution and OOD samples often hinges on 3D spatial properties, such as bond lengths or steric clashes, methods that do not account for the precise spatial positioning of nodes would fail to detect ”geometric outliers” that are topologically valid but physically impossible. Our work addresses this by combining graph features and 3D geometry into a single representation for the diffusion process.
Diffusion models for protein-ligand interactions
Score-based generative models and Denoising Diffusion Probabilistic Models (DDPMs) have achieved state-of-the-art results generating new protein-ligand interactions [17, 14, 29]. However, these works primarily focused on the generative task of sampling novel, stable conformations. The potential of using the probability flow ODE formulation [32] of a diffusion model for OOD detection in 3D molecular complexes remains unexplored. Our work bridges this gap by adapting the equivariant diffusion backbone from DiffSBDD [29] for precise OOD detection, thereby providing a much-needed OOD detection capability to the growing field of machine learning for protein-ligand interactions.
@@
Figure S1: Error-log-likelihood relation follows a scaled and shifted exponential: Scatterplot illustrating the relationship between the log-likelihoods assigned by the diffusion model (x-axis) and the corresponding GEMS binding affinity prediction errors (y-axis) across the validation, CASF2016, and six out-of-distribution (OOD) datasets. A hierarchical fitting of five scaled and shifted exponential curves shows that log-likelihoods provide a rough estimate of the expected GEMS prediction error. Across all non-training complexes ($N=6223$), 75.8% had predicted errors that fell within the bounds of the fitted exponential curves, with 8.84% high outliers (errors above the fit) and 15.39% low outliers (errors below the fit). To allow fair comparison across datasets with different label spreads, absolute errors are normalized by the label variance, preventing artificially low error values on narrowly distributed datasets. Exponential curves are fit to a maximum of 500 randomly sampled points from each non-training distribution. Only 100 randomly sampled complexes from each datasets are shown in the to improve visibility.
S1.3 Evaluation of sensitivity to ligand size and chemical features
@@
S1.4 Intra-Cluster Ligand Diversity
@@
Figure S3: Intra-cluster ligand diversity of all datasets sorted by cluster tightness in ascending order. Three robust intra-cluster diversity metrics are included: 1) Median distance to medoid measures typical cluster tightness by computing the median distance from the most representative point to all other points. 2) 90th percentile radius captures robust cluster size by identifying the distance that encompasses 90% of points from the medoid. 3) Pairwise inter-quartile mean (IQM) quantifies overall cluster cohesion by averaging distances between all unique pairs within the middle 50% of the pairwise distance distribution, providing resistance to outliers.
S1.5 Sampling Evaluation
@@
S1.6 PF-ODE likelihoods as predictors of predictive error
@@
S1.6.1 Relative entropy along the PF-ODE
@@
KL divergence and relative Fisher information
@@
$D_{t}$ $\;:=\;\mathrm{KL}\bigl(p_{t}\,\|\,q_{t}\bigr)\;=\;\int_{\mathbb{R}^{d}}p_{t}(x)\,\log\!\frac{p_{t}(x)}{q_{t}(x)}\,\mathrm{d}x,$ (S1.5)
$I_{t}$ $\;:=\;\int_{\mathbb{R}^{d}}p_{t}(x)\,\bigl\|\nabla_{x}\log p_{t}(x)-\nabla_{x}\log q_{t}(x)\bigr\|_{2}^{2}\,\mathrm{d}x.$ (S1.6)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}D_{t}\;=\;\int_{\mathbb{R}^{d}}p_{t}(x)\,\bigl(v^{\star}(x,t)-v_{\phi}(x,t)\bigr)^{\top}\Bigl[\nabla_{x}\log p_{t}(x)-\nabla_{x}\log q_{t}(x)\Bigr]\,\mathrm{d}x.$ (S1.7)